SLC,A fancy footwear manufacturing company has an obligation to pay MXN 14 million in 30 days for a recent shipment from Mexico.

The CFO of SLC is contemplating hedging the company’s MXN exposure on this transaction. He collects the below information regarding the interest rates and exchange rates, from her forex trader:

Spot Rate: MXN 20.08 / USD

Forward Rate: MXN 20.28 / USD

30-day Put Option on USD MXN 19.50 / USD: 1% Premium

30-day Call Option on USD MXN 20.50/ USD: 3% Premium

USD 30-day interest rate (annualized): 7.5%

MXN 30-day interest rate (annualized): 15%


You are required to answer the below questions to assist the CFO:-


a. What are the hedging options available to SLC? What is the hedged cost of SLC payable using a forward market hedge and using a put option hedge?

 

b. What is the hedged cost of SLC payable using money market hedge? 



1
Expert's answer
2021-09-13T16:54:27-0400

A) SCL can use Forward, Call option, and money market.

Forward cost:

"14000000/20.08—14000000/20.28=6875.85\ USD"

Options cost:

"14000000/20.08-14000000/19.5-14000000/19.5*0.01=-27917\ USD"

b) "14000000/(1+0.15*30/365)/20.08=688719\ USD"

"688719*(1+0.075*30/365)=692964\ USD"

"14000000/20.08-692964=4247\ USD"